### v1

# Edge-Spreading Raptor-Like LDPC Codes for 6G Wireless Systems

Abstract
Next-generation channel coding has stringent demands on throughput, energy consumption, and error rate performance while maintaining key features of 5G New Radio (NR) standard codes such as rate compatibility, which is a significant challenge. Due to excellent capacity-achieving performance, spatially-coupled low-density parity-check (SC-LDPC) codes are considered a promising candidate for next-generation channel coding. In this paper, we propose an SC-LDPC code family called edge-spreading Raptor-like (ESRL) codes for 6G. Unlike other SC-LDPC codes that adopt the structure of existing rate-compatible LDPC block codes before coupling, ESRL codes maximize the possible locations of edge placement and focus on constructing an optimal coupled matrix. Moreover, a new graph representation called the unified graph is introduced. This graph offers a global perspective on ESRL codes and identifies the optimal edge reallocation to optimize the spreading strategy. We conduct comprehensive comparisons of ESRL codes and 5G-NR LDPC codes. Simulation results demonstrate that when all decoding parameters and complexity are the same, ESRL codes have obvious advantages in error rate performance and throughput compared to 5G-NR LDPC codes, making them a promising solution towards next-generation channel coding.
Index Terms:
LDPC codes, SC-LDPC codes, rate-compatible, Raptor-like, wireless communications, 6G.
I Introduction
Next-generation channel coding targets a peak data rate of $200$ Gbps to $1$ Tbps with an energy efficiency target of around $1$ pJ/bit [1], representing 10×10$\times$10 × to 100×100$\times$100 × more stringent requirements compared to 5G New Radio (NR) [2]. As standards continue to evolve, channel coding also needs to meet diverse key performance indicators (KPIs), including higher throughput, enhanced reliability, lower power consumption, reduced computational complexity, rate compatibility, and new capabilities [3], leading to a significant challenge.
As a key technique in 5G-NR [4], low-density parity-check (LDPC) codes [5] have achieved significant success due to capacity-approaching performance [6] and parallel processing capabilities under belief propagation (BP) decoding [7]. For finite-length LDPC codes, though BP offers the advantage of reduced decoding complexity, the BP decoding threshold is sub-optimal compared to the maximum a-posteriori probability (MAP) threshold. Hence, spatially-coupled LDPC (SC-LDPC) codes (also called LDPC convolutional codes) were proposed in [8] and [9], by superimposing LDPC block codes with a convolutional structure. In general, block codes have a simple structure that is easy to implement in hardware, but they have increased complexity and latency when handling frames with large blocklengths. In contrast, convolutional codes are effective at decoding long frames but suffer from sequential processing of windowed decoding that introduces decoding latency. SC-LDPC codes, offering the best of both worlds, achieve the MAP threshold of the underlying block ensembles using BP decoding on a coupled ensemble [9, 10]. Besides, the compact and regular structure of underlying blocks also simplifies the decoding of long frames, and state-of-the-art SC-LDPC decoder implementations have reported high throughput in the range of hundreds of Gbps [11]. These ideas have attracted widespread interest for next-generation wireless and optical fiber communications [3, 12].
For wireless channels, rate compatibility and blocklength scalability are necessary features for channel coding. To address the time-varying nature of wireless channels, rate-compatible channel codes support incremental redundancy transmission, thereby maximizing the throughput under hybrid automatic repeat-request (HARQ) schemes [13]. Specifically, a rate-compatible code family is a nested sequence of codes with different rates, where higher rate codes are embedded into lower rate codes [14]. An initial codeword is transmitted at a high rate, and in case of erroneous reception or failed decoding, additional partial redundancy is sent until all redundancy of the mother code is used up. Numerous works on rate-compatible LDPC codes have been proposed in [14, 15, 16, 17] using puncturing and extension. Moreover, the design of LDPC codes can begin with the construction of a relatively small bipartite graph called the protograph [18]. A well-designed protograph applies a copy-and-permute operation (also called lifting) to obtain larger graphs with nodes, thus leading to longer-blocklength LDPC codes [19]. In this context, a protograph-based Raptor-like (PBRL) LDPC code family that meets the above requirements is a very popular one [17]. Sharing many similarities with LDPC codes, the introduction of Luby transform (LT) codes [20] and Raptor codes [21], which can virtually generate any desired rate, form a novel coding paradigm of rateless codes. Subsequently, PBRL codes restrict the protograph to a Raptor-like structure, achieving both rate compatibility and blocklength adjustment simultaneously. Hence, PBRL LDPC codes have been adopted into 5G-NR successfully [4]. However, the complexity of BP decoding and hardware increases as the blocklength grows in 5G-NR. As indicated by the state-of-the-art 5G-NR LDPC decoder implementation [22], the memory overhead required to accommodate the maximum blocklength occupies a significant part of the decoder core area (over $60\%$) and results in complex routing. This limits peak throughput to around $20$ Gbps for 5G-NR LDPC codes, far away from stringent demands beyond 5G.
Given the advantages of SC-LDPC codes over traditional LDPC block codes, developing a rate-compatible SC-LDPC code family has emerged as a promising solution towards next-generation channel coding. Note that SC-LDPC codes can be constructed from LDPC block codes using an edge-spreading technique [23]. Several SC-LDPC code constructions tailored to a specific rate were presented in [24, 25, 26, 27, 28, 29], where most of them focus on optimizing edge spreading starting from a given LDPC block code, mitigating the emergence of harmful graph structures. For instance, [26] proposed a heuristic search combined with a check-and-flip process to construct short-cycle-free SC-LDPC codes by generating multiple auxiliary matrices. However, the pseudo-random nature of [26] leads to significant complexity, particularly as the matrix size and coupling width increase. To address this issue, an edge-spreading matrix [30] has been employed in recent works [25, 27, 28, 29] to provide a more compact representation of edge spreading. This enables a simple method to identify harmful structures in SC-LDPC codes, regardless of the increase in coupling width. Sima et al. [25] imposed constraints on the edge-spreading matrix to reduce dominant trapping sets and avoid error floors, and [28] analyzed the relationship between short cycles and edge spreading to design short-cycle-free SC-LDPC codes with minimal coupling width. To further incorporate rate compatibility [31, 32, 33, 34, 35], Si et al. introduced the first rate-compatible SC-LDPC code to achieve the capacity of the binary erasure channel (BEC) [31]. Zhou et al. designed a family of robust rate-compatible punctured SC-LDPC codes by ensuring the recoverability of punctured bits and avoiding harmful graph structures. Protograph-based rate-compatible SC-LDPC codes were proposed in [33, 34, 35] following a simple idea. Namely, a conventional PBRL protograph is first designed, and then the edge-spreading technique is directly applied to construct a coupled code, with both procedures being independent. Notably, [35] achieved performance gains by simply splitting the 5G-NR LDPC matrix as the foundation for SC-LDPC codes.
However, these SC-LDPC works either focus on optimizing code constructions tailored to a specific rate [24, 25, 26, 27, 28, 29] or they apply the structure of existing rate-compatible LDPC block codes to generate SC-LDPC codes without considering the influence of the coupling on performance [33, 34, 35], thus not exploring the full potential of SC-LDPC codes. Consequently, there is a need for a holistic design method to construct SC-LDPC codes that simultaneously meet diverse demands, including rate compatibility, protograph design with coupling, threshold analysis, removal of harmful graph structures, and high throughput potential. Moreover, it is necessary to conduct comprehensive comparisons of the designed SC-LDPC codes and standard LDPC codes, especially when all decoding parameters and complexity are consistent. This would provide a clear assessment of the potential of SC-LDPC codes for next-generation channel coding.
Contributions: This paper proposes an SC-LDPC code family called edge-spreading Raptor-like (ESRL) LDPC codes. Unlike PBRL codes [17], ESRL codes maximize the possible locations of edge placement before coupling and focus on constructing an optimal coupled matrix. Hence, ESRL codes show outstanding performance while maintaining extensive rate compatibility. The structure, properties, and code profile of ESRL codes are discussed in detail.
Moreover, we propose a novel graph representation called the unified graph. Unlike other edge-spreading optimizations involved with complex processing [27, 28, 29, 30], this graph innovatively offers a global perspective on ESRL codes, identifying the optimal edge reallocation to optimize the spreading strategy. We also introduce modified coupled reciprocal channel approximation (RCA) to analyze the decoding threshold of ESRL codes, demonstrating that ESRL codes enjoy better thresholds across a wide range of rates. Based on these tools, a complete design flow of ESRL codes is presented.
This paper also proposes a high-throughput decoding algorithm tailored to ESRL codes, called semi-layered multi-engine (SLME) decoding. Based on features of ESRL codes, SLME decoding decomposes the original BP decoding into multiple identical sub-decoders that operate in parallel in a more compact manner compared to [12]. This approach significantly enhances the decoding parallelism for ESRL codes.
Finally, we provide a design example of ESRL codes and conduct comprehensive comparisons with 5G-NR LDPC codes [4]. As summarized in Fig. 1, ESRL codes preserve several advantages and properties of 5G-NR LDPC codes while further enhancing error rate performance and throughput. When all decoding parameters and complexity are the same, at the rate of $0.52$, ESRL codes outperform 5G-NR LDPC codes by $0.95$ dB at five iterations or enhance peak throughput by $50.1\%$ using a similar hardware architecture [22].
Figure 1: Comparisons of ESRL codes and 5G-NR LDPC codes [4]. More details are discussed in Section VI-B.
Outline: The remainder of this paper is organized as follows. Section II introduces the symbol definitions and reviews in compact form the preliminaries of LDPC codes. Section III introduces the structure, properties, and code profile of ESRL codes. Section IV outlines the design and optimization procedures for ESRL codes, including the unified graph and modified coupled RCA. In Section V, we present the puncturing scheme and SLME decoding for ESRL codes. Section VI provides a design example of ESRL codes and compares it with 5G-NR LDPC codes. Finally, Section VII concludes this paper.
II Preliminaries
In this paper, we use the following definitions. Boldface lowercase letters $\bm{u}$ denote vectors, where $\bm{u}_{i}$ refers to the $i$-th element of $\bm{u}$ and $\bm{u}_{i:j}$ is the sub-vector $[\bm{u}_{i},\bm{u}_{i+1},\dots,\bm{u}_{j}]$$i\leq j$, and the null vector otherwise. Boldface uppercase letters $\mathbf{B}$ represent matrices, with $\mathbf{B}_{i,j}$ denoting the element in the $i$-th row of the $j$-th column. Blackboard letters, such as $\mathbb{S}\triangleq\{\cdot\}$, denote sets with $|\mathbb{S}|$ being the cardinality.
LDPC block codes are specified by a sparse $M\times N$ parity-check matrix (PCM) $\mathbf{H}$, which can be further described by a bipartite Tanner graph [36] with the information length of $K=N-M$. If $\mathbf{H}_{c,v}=1$ for $0\leq c<M$ and $0\leq v<N$, the $c$-th check node (CN) is connected to the $v$-th variable node (VN) on the Tanner graph. We use $\mathbb{V}_{c}$ to represent the set of adjacent VNs for the $c$-th CN and we use $\mathbb{C}_{v}$ to denote the set of neighbors for the $v$-th VN. Let $|\mathbb{V}_{c}|=d_{c}$ and $|\mathbb{C}_{v}|=d_{v}$ refer to the row degree of the $c$-th CN and the column degree of the $v$-th VN, respectively. Due to the quasi-cyclic (QC) property [19], a QC-LDPC code can be described by a more structured $m\times n$ protomatrix $\mathbf{B}$ with the information length of $k=n-m$. where each entry $\mathbf{B}_{c,v}$ can be extended by a $Z\times Z$ cyclic-shifting identity matrix, with $Z$ being the lifting size. To focus on the protomatrix design in this paper, $\mathbf{B}_{c,v}$ denotes only the number of connected edges between the $c$-th CN and the $v$-th VN instead of a specific cyclic shifting value. Similarly, $\mathbf{B}$ can be transformed into a protograph $\mathcal{G}$ [18], defined as $\mathcal{G}\triangleq\{\mathbb{V}\cup\mathbb{C},\mathbb{E}\}$, where $\mathbb{E}$ is the set of all edges in $\mathcal{G}$. Let $e_{c}^{v}$ denote the directed edge from the $c$-th CN to the $v$-th VN, and $e_{v}^{c}$ is the same edge with the opposite direction. Note that all graph concepts are based on protographs throughout this paper and we omit the term ‘QC’ for simplicity.
Through a convolutional structure, a chain of multiple identical protographs can be organized to derive a protograph-based SC-LDPC code with a coupling width of $\omega$. To this end, after edge spreading at each spatial position, $\mathbf{B}$ is split into $\omega+1$ sub-matrices $\mathbf{B}_{i}$$i\in[0,\omega]$, i.e., $\mathbf{B}=\sum_{i=0}^{\omega}\mathbf{B}_{i}$, each edge in the original $\mathbf{B}$ is assigned to exactly one of the sub-matrices. To construct a protomatrix $\mathbf{B}_{\text{SC}}$ of a finite-length SC-LDPC code, $L$ replicas of a series of sub-matrices $\mathbf{B}_{i}$$i\in[0,\omega]$, are coupled to generate $\mathbf{B}_{\text{SC}}$ and the corresponding protograph $\mathcal{G}_{\text{SC}}$, where $L$ is the coupling length. Consequently, a PCM $\mathbf{H}_{\text{SC}}$ can also be derived from a series of sub-PCMs $\mathbf{H}_{i}$$i\in[0,\omega]$ with a lifting size $Z$. Let $\bm{c}_{[0,L-1]}\triangleq[\bm{c}_{0},\bm{c}_{1},\dots,\bm{c}_{L-1}]$ denote a codeword of the SC-LDPC code, and $\bm{c}_{[i]}\triangleq[\bm{u}_{[i]},\bm{r}_{[i]}]$ is the codeword of the $i$-th block (also called batch) in an HARQ process, where $\bm{u}_{[i]}$ is the information vector and $\bm{r}_{[i]}$ is the corresponding parity-check vector. Note that the above $\mathbf{B}_{\text{SC}}$, $\mathcal{G}_{\text{SC}}$, $\mathbf{H}_{\text{SC}}$, and $\bm{c}_{[0,L-1]}$ temporarily ignore the last redundancy tail, which we will explain in detail below.
III Edge-Spreading Raptor-Like LDPC Codes
In this section, we introduce the structure, properties, and code profile of the proposed ESRL code family.
Figure 2: Uncoupled protograph and protomatrix for an ESRL code with a HRC, an IRC, and an SSC. Upper right corner are empty to ensure rate compatibility.
III-A Structure of ESRL Codes
Fig. 2 illustrates the structures of the uncoupled protograph $\mathcal{G}$ and uncoupled protomatrix $\mathbf{B}$ of an ESRL code, which consists of three components: a highest-rate code (HRC), an incremental redundancy code (IRC), and a supplementary-spreading code (SSC). Unlike conventional protographs of PBRL codes (such as the ones used in 5G-NR) that feature a rightmost identity matrix [17], an SSC matrix is used to maximize the possible locations for edge placement. Let (1) denote $\mathbf{B}$ of an ESRL code
$\mathbf{B}=\left[\begin{array}[]{c c}\mathbf{B}_{\text{HRC}}&\mathbf{0}\\ \mathbf{B}_{\text{IRC}}&\mathbf{B}_{\text{SSC}}\\ \end{array}\right].$ (1)
Previous approaches to designing rate-compatible SC-LDPC codes [31, 32, 33, 34, 35] typically design a good uncoupled protomatrix first, followed by coupling. For instance, [33, 34, 35] attempted to generate rate-compatible SC-LDPC codes by directly applying edge spreading to a PBRL protograph. However, as shown in Fig. 3, decoding of SC-LDPC codes now occurs on the coupled $\mathbf{B}_{\text{SC}}$ rather than the uncoupled $\mathbf{B}$ like in LDPC block codes. This highlights a fundamental difference of our ESRL codes compared to [31, 32, 33, 34, 35], i.e., our target is to design an uncoupled protomatrix $\mathbf{B}$ that ultimately results in a good coupled protomatrix $\mathbf{B}_{\text{SC}}$ after edge spreading, regardless of whether $\mathbf{B}$ by itself is optimal. Given the Raptor-like structure inherent to ESRL codes, establishing a good Raptor-like structure after edge spreading is clearly more important than good performance before coupling. Define a partially coupled chain as
$\mathbf{U}=\left[\mathbf{B}_{\omega},\dots,\mathbf{B}_{1},\mathbf{B}_{0}\right].$ (2)
Given the convolutional structure of SC-LDPC codes, decoding can be further decomposed into multiple identical engines (shown in Fig. 3) working on different $\mathbf{U}$ blocks at various spatial positions in parallel (also called multiple-engine decoding in [12]). Intuitively, decoding on $\mathbf{U}$ largely determines the overall performance of SC-LDPC decoding. However, a conventional PBRL structure restricts the design space to find the optimal $\mathbf{B}_{\text{SC}}$, since its rightmost SSC is only an identity matrix and must be fully allocated to $\mathbf{B}_{0}$ for encoding [37]. This reduces the possible locations for edge placement and leaves the right parts of $\mathbf{B}_{\omega},\mathbf{B}_{\omega-1},\dots,\mathbf{B}_{1}$ in (2) empty. The proportion of empty parts increases to $67.6\%$ across medium to low rates, leading to performance loss [33, 34, 35]. Hence, SC-LDPC codes need a new structure for the uncoupled protomatrix to enhance message propagation on $\mathbf{U}$. Given that [14, 15, 16] have indicated that all rate-compatible LDPC codes must adhere to a lower triangular constraint on the right side, either uncoupled or coupled, the ESRL code also follows this constraint to preserve the upper right corner of (1) empty and keep the SSC lower triangular. Note that despite the motivation of ESRL codes to design the optimal $\mathbf{B}_{\text{SC}}$ after edge spreading, the ESRL code still retains the fundamental HRC and IRC parts of traditional PBRL codes. This means an ESRL decoder can be naturally compatible with PBRL codes of appropriate sizes such as 5G-NR LDPC codes.
The ESRL structure also resembles Raptor codes [21], but with key differences. First, while the HRC is identical to the precode in a Raptor code, the ESRL code is systematic, with the codeword of each batch represented as $\bm{c}_{[i]}=[\bm{u}_{[i]},\bm{r}_{[i]}],0\leq i<L$. This design allows explicit transmission of the HRC VNs in all batches and supports decoding at the highest rate, which is infeasible in a Raptor code. Second, similar to the LT component in Raptor codes, XOR operations in the ESRL code generate the VNs of the IRC and the SSC, producing additional parity bits row-wise. However, these connections in the ESRL code are pre-defined rather than arbitrary. Third, compared to two-phase BP decoding of Raptor codes or joint BP decoding on PBRL codes, the ESRL code can support a more flexible windowed BP decoding to slide over the coupled protograph $\mathcal{G}_{\text{SC}}$ in a low-complexity manner [38]. More details on decoding are discussed in Section V-B.
III-B Properties of ESRL Codes
In this section, we present some important properties of ESRL codes. First, the ESRL code is a time-invariant SC-LDPC code with no parallel edges, i.e., each entry of $\mathbf{B}$ is binary, and no non-zero entries in $\mathbf{B}_{i},\forall i\in[0,\omega]$ occupy the same position. The time-invariant nature and the absence of parallel edges ensure that the ESRL code can be decomposed into multiple independent identical engines that decode on different $\mathbf{U}$ blocks in parallel, without causing memory access conflicts (assuming the decoding schedule is consistent at each engine). This results in ESRL codes that allow for significant decoding parallelism to potentially achieve the high throughput demand of next-generation channel coding, as demonstrated in Section V-B and Section VI. It is noteworthy that although parallel edges can reduce the decoding threshold [17, 39] and can be easily assigned to $\omega$ sub-matrices in SC-LDPC codes without the need for two-step lifting [39], this performance gain sacrifices parallelism in an ESRL decoder. The absence of parallel edges, as mentioned in [29], can also simplify the theoretical tools described in Section IV to optimize ESRL codes.
Figure 3: The code profile of the ESRL code, including an uncoupled protomatrix $\mathbf{B}$, an edge-spreading matrix $\mathbf{T}$, and a tail matrix $\mathbf{Q}$.
The HRC protograph is the protograph design of a good code at a given rate, which has been well studied in [39]. Following the example of [39], we set the minimum column degree to at least $3$ to ensure linear minimum distance growth with increasing blocklength. Besides, a punctured VN that is fully connected to all CNs in the HRC is placed in the first column to reduce the decoding threshold. The IRC component aims to provide reliability back to the VNs of the original HRC. In a complete ESRL matrix with the lowest rate, the CNs of the IRC must eventually connect back to all VNs of the HRC. Following [40], each row of the IRC has at least two edges to retain linear minimum distance growth with increasing blocklength. Besides, the punctured VN of the precode connects to all CNs in the IRC. Unlike the use of parallel edges in [17], the HRC and the IRC of ESRL codes do not involve any parallel edges.
The SSC protograph is to enhance message propagation on $\mathbf{U}$, especially for medium to low rates. In the SSC, all diagonal elements are set to ‘$1$’s to facilitate encoding. Note that the SSC is connected to and primarily improves the reliability of VNs associated with redundancy, which are generally less critical than the VNs in the HRC, thus the weight of the SSC should be kept lower than that of the IRC. While the rate reduces, we increase the SSC weight to avoid the right side being too sparse, but by experimental trial, this proportion is at most $30\%$ of the IRC weight.
III-C Code Profile of ESRL Codes
In this section, we present the code profile of the ESRL code family. Compared with LDPC block codes, SC-LDPC codes contain more information, such as a coupling relationship, which cannot be described by a single uncoupled protomatrix $\mathbf{B}$. Although a straightforward representation using the coupled protomatrix $\mathbf{B}_{\text{SC}}$ is feasible, such a description would contain a significant amount of redundant information due to the presence of $L$ replicas of $\mathbf{U}$. In this paper, we provide a simpler method to represent ESRL codes. Namely, a complete ESRL code can be described using three matrices: an uncoupled protomatrix $\mathbf{B}$, an edge-spreading matrix $\mathbf{T}$, and a tail matrix $\mathbf{Q}$, as illustrated in Fig. 3(c).
The uncoupled protomatrix $\mathbf{B}$ outlines the parent structure of ESRL codes, as explained in Section III-A. To convert $\mathbf{B}$ to $\mathbf{B}_{\text{SC}}$, following [30, 29], we introduce an edge-spreading matrix $\mathbf{T}$, instead of explicitly recording $\omega+1$ sub-matrices as in [23, 26]. The matrix $\mathbf{T}$ of size $m\times n$ determines how the edges of the uncoupled protograph $\mathcal{G}$ are spread to the coupled protograph $\mathcal{G}_{\text{SC}}$. Specifically, the elements of $\mathbf{T}$ indicate to which sub-matrices the edges of $\mathbf{B}$ are assigned. We set $\mathbf{T}_{c,v}=-1$ if $\mathbf{B}_{c,v}=0$. Otherwise, $\mathbf{T}_{c,v}=i$, $i\in[0,\omega]$, implies that the entry at row $c$ and column $v$ of $\mathbf{B}_{c,v}$ is assigned to the corresponding position in $\mathbf{B}_{i}$. Let $\delta_{i}(\mathbf{T}_{c,v})$ be a function that returns a boolean value of ‘$1$’ if $\mathbf{T}_{c,v}$ equals $i$, and ‘00’ otherwise. As a result, we can generate $\omega+1$ sub-matrices $\mathbf{B}_{0},\mathbf{B}_{1},\dots,\mathbf{B}_{\omega}$ explicitly as
$\mathbf{B}_{i}\!\!=\!\!\left[\!\!\!\begin{array}[]{c c c}\mathbf{B}_{0,0}\cdot\delta_{i}(\mathbf{T}_{0,0})&\cdots&\mathbf{B}_{0,n-1}\cdot\delta_{i}(\mathbf{T}_{0,n-1})\\ \vdots&\ddots&\vdots\\ \mathbf{B}_{m\!-\!1,0}\!\cdot\!\delta_{i}(\mathbf{T}_{m\!-\!1,0})\!\!\!\!\!&\cdots&\!\!\!\!\mathbf{B}_{m\!-\!1,n\!-\!1}\!\cdot\!\delta_{i}(\mathbf{T}_{m\!-\!1,n\!-\!1})\\ \end{array}\!\!\!\right].$ (3)
Furthermore, the termination of finite-length SC-LDPC codes results in an unavoidable rate loss at the boundary. To mitigate this issue, one can either append a redundancy matrix $\mathbf{Q}$ at the end [41, 37] or reduce the information length of the last block [42] to satisfy encoding. ESRL codes adopt the former approach, as it has a lower rate loss. Previous works generally use a simple identity matrix for $\mathbf{Q}$ [41, 37], but this leads to serious error floors at medium to low rates. In contrast, we adopt a tri-diagonal matrix for $\mathbf{Q}$ with a lower triangular structure to ensure that most column degrees are either $2$ or $3$. For rate compatibility, the matrix $\mathbf{Q}$ must remain full-rank even after pruning any $j$ rows and $j$ columns from the end. Based on the three matrices $\mathbf{B}$, $\mathbf{T}$, and $\mathbf{Q}$, we explicitly define the coupled protomatrix $\mathbf{B}_{\text{SC}}$ of the ESRL code as follows
$\mathbf{B}_{\text{SC}}\!=\!\left[\!\!\begin{array}[]{ccccccc}\mathbf{B}_{0}&&&&&\cdots&\mathbf{0}\\ \mathbf{B}_{1}&\mathbf{B}_{0}&&&&&\vdots\\ \vdots&\mathbf{B}_{1}&\ddots&&&&\\ \mathbf{B}_{\omega}&\vdots&\ddots&\mathbf{B}_{0}&&&\\ &\mathbf{B}_{\omega}&\ddots&\mathbf{B}_{1}&\mathbf{Q}_{1}&\cdots&\mathbf{0}\\ \vdots&&\ddots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\cdots&&\mathbf{B}_{\omega}&\mathbf{0}&\cdots&\mathbf{Q}_{\omega}\\ \end{array}\!\!\right]\!.$ (4)
Without loss of generality, we decompose the tail matrix $\mathbf{Q}$ in (4) into a series of matrices $\mathbf{Q}_{1},\dots,\mathbf{Q}_{\omega}$ (each of size $m\times m$) on the diagonal, where an example of $\mathbf{Q}$ is illustrated in Fig. 9. It is noteworthy that the three matrices $\{\mathbf{B},\mathbf{T},\mathbf{Q}\}$ allow for flexible pruning from the end, enabling the ESRL code to support a wide range of rates with one code matrix. Let $\rho$ denote the number of punctured VNs in $\mathbf{B}$, we can define the code rate $R$ of the ESRL code as
$R=\frac{k\cdot L}{n\cdot L+m\cdot\omega-\rho\cdot(L+\omega)}.$ (5)
IV Design of ESRL Codes
In this section, we outline the procedures for the design and optimization of a good ESRL code in detail.
IV-A Unified Graph for ESRL Codes
Inspired by the aforementioned code profile, we propose a novel graph representation for ESRL codes, which we call the unified graph $\mathcal{G}_{\text{U}}$. This graph contains all essential code information and enables a comprehensive illustration of all possible graph structures within an SC-LDPC protograph. Unlike traditional approaches, a single unified graph $\mathcal{G}_{\text{U}}$ can integrate the relationships and dependencies among the matrices $\mathbf{B}$ and $\mathbf{T}$ in ESRL codes, [^1] Given that the tail matrix $\mathbf{Q}$ has a fixed structure and is not involved in edge spreading, we ignore the presence of $\mathbf{Q}$ to simplify the optimization on $\mathcal{G}_{\text{U}}$. Note that $\mathcal{G}_{\text{U}}$ is a protograph concept. thereby offering a more global view of the code construction. The construction of $\mathcal{G}_{\text{U}}$ is built on the bipartite graph derived from $\mathbf{B}$, where edges are annotated by different line types based on the corresponding elements of $\mathbf{T}$. For example, Fig. 4 depicts the unified graph $\mathcal{G}_{\text{U}}$ derived from $\mathbf{B}$ and $\mathbf{T}$ with $\omega=1$, where dashed and solid lines distinguish whether edge $e$ is assigned to $\mathbf{B}_{0}$ or $\mathbf{B}_{1}$, respectively. In this section, we present a message-passing algorithm on $\mathcal{G}_{\text{U}}$ to count and eliminate short cycles of $\mathcal{G}_{\text{SC}}$ efficiently, which inspires a good spreading strategy on $\mathcal{G}_{\text{U}}$ to enhance the performance of ESRL codes.
IV-A1 Cycle Counting over the Unified Graph
We first define cycles of length $l$ (referred to as $l$-cycles) in a protograph $\mathcal{G}$. An $l$-cycle corresponding to a sequence of $l/2$ VNs and $l/2$ CNs in (6), or equivalently, a sequence of $l$ edges in (7), starts and ends at the same node.
$\{v_{0},c_{0},v_{1},c_{1},\ldots,v_{l/2-1},c_{l/2-1},v_{0}\}.$ (6)
$\{e_{v_{0}}^{c_{0}},e_{c_{0}}^{v_{1}},e_{v_{1}}^{c_{1}},e_{c_{1}}^{v_{2}},\dots,e_{v_{l/2-1}}^{c_{l/2-1}},e_{c_{l/2-1}}^{v_{0}}\}.$ (7)
For rigour, we only consider simple cycles in this paper, i.e., tailless backtrackless closed (TBC) walks in a graph. After edge spreading from $\mathcal{G}$ to $\mathcal{G}_{\text{SC}}$, we can get Remark 1.
Remark 1.
The number of $l$-cycles in $\mathcal{G}$ is the upper bound of the number of $l$-cycles in $\mathcal{G}_{\text{SC}}$ if repeated cycles in $L$ replicas are counted only once.
Proof.
Let $\phi$ denote the accumulated path value for an $l$-cycle, as calculated by
$\phi=\Delta\mathbf{T}_{0}+\Delta\mathbf{T}_{1}+\Delta\mathbf{T}_{2}+\ldots+\Delta\mathbf{T}_{l/2-1},$ (8)
where $\Delta\mathbf{T}_{i}=\mathbf{T}_{c_{i},v_{i}}-\mathbf{T}_{c_{i},v_{\text{mod}(i+1,l/2)}}$. Note that $\phi$ tracks the path of how an $l$-cycle passes through $\omega+1$ sub-matrices in SC-LDPC codes. Since an $l$-cycle in $\mathcal{G}$ remains in $\mathcal{G}_{\text{SC}}$ only if $\phi=0$ [30, 29] otherwise broken, $\mathcal{G}_{\text{SC}}$ will not generate any new unique cycle compared to $\mathcal{G}$. ∎
Figure 4: A unified graph $\mathcal{G}_{\text{U}}$ example, where dashed black lines and solid black lines are assigned to $\mathbf{B}_{0}$ and $\mathbf{B}_{1}$, respectively.
Subsequently, we introduce a message-passing algorithm over $\mathcal{G}_{\text{U}}$ to efficiently count cycles for ESRL codes. In general, a message-passing algorithm computes messages at the nodes and propagates them along the edges of a graph, where the calculation of the message sent along an edge $e$ excludes the message previously received along $e$ (refers to this property as extrinsic[7]. Let $\gamma$ denote a basic message, and an example of an extrinsic message-passing process is illustrated in Fig. 4. Hence, the messages over the edges of $\mathcal{G}_{\text{U}}$ are all monomials of $\gamma$ and the message-passing operation is multiplication. Besides, we need to record the extra accumulated path value $\phi$ of each $\gamma$ over $\mathcal{G}_{\text{U}}$, due to using edge spreading. Let $t$ denote the index of discrete time, and we call each $t$ as one step, corresponding to either a VN-to-CN or CN-to-VN phase, where $t$ always starts from 00 of the VN-to-CN phase. We denote $\mathcal{M}_{v,c}^{(t)}$ as the VN-to-CN message along $e_{v}^{c}$ and $\mathcal{M}_{c,v}^{(t)}$ as the CN-to-VN message along $e_{c}^{v}$. The update rules are defined in (9) and (10), and the $\phi$ of each message is also tracked dynamically to verify whether the cycle remains or not in $\mathcal{G}_{\text{SC}}$ based on Remark 1.
Herein, we present how to use $\mathcal{G}_{\text{U}}$ count cycles that include a specific VN. For the $\bar{v}$-th VN, the initial message $\mathcal{M}_{\bar{v},c}^{(0)}$ at $t=0$ along the edge $e_{\bar{v}}^{c}$ is denoted by $\gamma_{c}^{\phi_{c}}$, where $\phi_{c}$ is initialized as $\mathbf{T}_{c,\bar{v}}$. We can summarize the VN-to-CN message-passing rules as
$\mathcal{M}_{v,c}^{(t)}=\prod\limits_{c^{\prime}\in\mathbb{C}_{v}\backslash c}h^{+}\left(\mathcal{M}_{c^{\prime},v}^{(t-1)},\mathbf{T}_{c^{\prime},v}\right),\quad t\geq 1.$ (9)
Similarly, the CN-to-VN message-passing follows
$\mathcal{M}_{c,v}^{(t)}=\prod\limits_{v^{\prime}\in\mathbb{V}_{c}\backslash v}h^{-}\left(\mathcal{M}_{v^{\prime},c}^{(t-1)},\mathbf{T}_{c,v^{\prime}}\right),\quad t\geq 1.$ (10)
The $h^{+}(\cdot)$ and $h^{-}(\cdot)$ functions are defined as (11) and (12) to update the accumulated path value $\phi$ in the current message. Let $\Phi(\cdot)$ denote a function to extract the $\phi$ from all $\gamma$ variables.
$h^{+}(\mathcal{M}_{c,v},\mathbf{T}_{c,v})=\Phi(\mathcal{M}_{c,v})+\mathbf{T}_{c,v}.$ (11)
$h^{-}(\mathcal{M}_{v,c},\mathbf{T}_{c,v})=\Phi(\mathcal{M}_{v,c})-\mathbf{T}_{c,v}.$ (12)
Notably, the general message format $\mathcal{M}$ that propagates along the edges in $\mathcal{G}_{\text{U}}$ is the product of multiple $\gamma$ variables, each labeled by individual path value $\phi$, where both the number and type of $\gamma$ are dynamic. At the $t\!=\!l-1$-th step, we will obtain the number of $l$-cycles that include the $\bar{v}$-th VN as
$Q_{\bar{v}}^{l}=f\left(p_{\bar{v}}^{l})\right/2,$ (13)
where the division by $2$ arises from the directed edges, as each cycle is counted twice. The function $f\left(p_{\bar{v}}^{l}\right)$ computes the number of all $\gamma$ variables with $\phi=0$ in the polynomial $p_{\bar{v}}^{l}=\sum_{c\in\mathbb{C}_{\bar{v}}}\mathcal{M}_{c,\bar{v}}^{(l-1)}$, i.e., only sequences with $\phi=0$ will remain as cycles in $\mathcal{G}_{\text{SC}}$ after edge spreading. Note that due to the extrinsic property mentioned before, the component formed by the initial message $\gamma_{c}$ sent along the edge $e_{\bar{v}}^{c}$ should be excluded from $\mathcal{M}_{c,\bar{v}}^{(l-1)}$ in (13[43]. Consequently, we repeat the above procedures for all VNs, $v\in\mathbb{V}$. The total number of $l$-cycles in the entire $\mathcal{G}_{\text{SC}}$ can be obtained as (14), where the accumulated result in (14) needs to be divided by $l$ as every $l$-cycle is counted $l$-times. Define the shortest cycle length in a graph as the girth $g$. The above approach can efficiently count $l$-cycles when $l\in\{4,6,\dots,2g-2\}$.
$Q^{l}=\left(\sum\limits_{v\in\mathbb{V}}Q_{v}^{l}\right)/\left(l/2\right).$ (14)
Example 1.
We provide an example in Fig. 5, illustrating the use of the message-passing algorithm over the unified graph $\mathcal{G}_{\text{U}}$ to count $4$-cycles and $6$-cycles containing $v_{4}$.
  1. 1.
    At $t=0$, $v_{4}$ sends the initial messages $\gamma_{1}^{1}$$\gamma_{2}^{0}$, and $\gamma_{3}^{0}$ along the adjacent edges $e_{v_{4}}^{c_{1}}$$e_{v_{4}}^{c_{2}}$, and $e_{v_{4}}^{c_{3}}$, respectively, where the initial path values are according to the corresponding elements of $\mathbf{T}$. All the messages sent by $v_{0}$, $v_{1}$, $v_{2}$, and $v_{3}$ along their edges are equal to $1$ and omitted. At $t=1$, all valid messages are updated based on (10) and (12) for the CN-to-VN phase.
  2. 2.
    At $t=2$, all VNs have valid incoming messages, which then pass extrinsic messages using multiplication of (9) and update $\phi$ by (11). At $t=3$, based on (13) we obtain the number of $4$-cycles as $Q_{v_{4}}^{4}=f\left(\gamma_{2}^{-1}+\gamma_{1}^{1}\gamma_{3}^{1}+\gamma_{2}^{-1}\right)/2=(0+0+0)/2=0$.
  3. 3.
    At $t=4$, all VNs propagate messages to their neighbors. For simplicity, we omit the messages that fail to return to $v_{4}$ at $t=5$ in the graph. Note that $t=5$, $\gamma_{2}^{-2}\gamma_{2}^{2}$ needs to be removed from $\mathcal{M}_{c_{2},v_{4}}^{(5)}$. Since the $\gamma_{2}$ variable is related to the initial message that passes along an edge and finally returns to $v_{4}$ by the same edge, it will not indicate a valid cycle. Therefore, we have $Q_{v_{4}}^{6}=f\left(p_{v_{4}}^{6}\right)=f\left(\gamma_{2}^{1}\gamma_{3}^{-1}\gamma_{3}^{0}+\gamma_{1}^{-1}\gamma_{3}^{-2}\gamma_{3}^{0}+\gamma_{1}^{0}\gamma_{1}^{1}\gamma_{2}^{0}\gamma_{2}^{2}\right)/2=(1+1+2)/2=2$.
Hence, this example illustrates that there are no $4$-cycles but two $6$-cycles passing through $v_{4}$ in $\mathcal{G}_{\text{U}}$.
IV-A2 Spreading Optimization over the Unified Graph
In this section, we present how to design a good edge-spreading strategy on the unified graph $\mathcal{G}_{\text{U}}$. When we have utilized $\mathcal{G}_{\text{U}}$ to explicitly count the cycles associated with each VN and the total number of cycles in $\mathcal{G}_{\text{SC}}$, $\mathcal{G}_{\text{U}}$ can further guide us in optimizing the spreading matrix to minimize short cycles.
Figure 5: An example of cycle counting on a unified graph $\mathcal{G}_{\text{U}}$.
Unlike [25, 24], there is no need to repeatedly run the above message-passing algorithm on $\mathcal{G}_{\text{U}}$ to update the cycle distribution when changes occur in $\mathcal{G}_{\text{U}}$. Instead, we can simply adjust the polynomial $p_{v}$ in (13) to measure the impact of reallocating a specific edge on the cycle distribution. It is noteworthy that the extra complexity of such an optimization on $\mathcal{G}_{\text{U}}$ is negligible. Specifically, if the edge $e_{v}^{c}$ is re-assigned from $\mathbf{T}_{c,v}$ to ${\widetilde{\mathbf{T}}_{c,v}}$ (where ${\widetilde{\mathbf{T}}_{c,v}}\in[0,\omega]$ and ${\widetilde{\mathbf{T}}_{c,v}}\neq\mathbf{T}_{c,v}$), let $\theta={\widetilde{\mathbf{T}}_{c,v}}-\mathbf{T}_{c,v}$, we modify the original $p_{v}$ by two steps: 1 Add $\theta$ to all accumulated path values $\phi$ of the initial $\gamma_{c}$ variable in the polynomial $p_{v}$ 2 Subtract $\theta$ from the $\phi$ of all $\gamma$ variables in the monomial passed to the VN $v$ along the edge $e_{c}^{v}$ (as both $e_{v}^{c}$ and $e_{c}^{v}$ belong to the same undirected edge, they are equivalent in the spreading process).
Following the example of Fig. 5, we observe that $v_{4}$ is involved in no $4$-cycles, but only two $6$-cycles after edge spreading. To eliminate these $6$-cycles, assuming only one edge allocation can be changed at a time, the adjustable edge candidate set is $\{e_{v_{4}}^{c_{1}},e_{v_{4}}^{c_{2}},e_{v_{4}}^{c_{3}}\}$ [^2] For simplicity, we only consider the edges with the VN-to-CN direction. and the corresponding initial spreading value set is $\{1,0,0\}$. We modify each spreading value in turn, thus resulting in polynomials $\{{\widetilde{p}_{\widetilde{\mathbf{T}}_{c_{1},v_{4}}}^{6}},{\widetilde{p}_{\widetilde{\mathbf{T}}_{c_{2},v_{4}}}^{6}},{\widetilde{p}_{\widetilde{\mathbf{T}}_{c_{3},v_{4}}}^{6}}\}$ as in
[…]
The difference between the original $p_{v_{4}}^{6}$ and the updated polynomials is highlighted in red. Hence, based on (15), we can quickly compute the number of $6$-cycles containing $v_{4}$ that can be removed as $\{0,1,1\}$. Note that despite one $6$-cycle removed by the modification of $\mathbf{T}_{2,4}$ from 00 to $1$, it will inevitably lead to two new $4$-cycles, i.e., $f\left(\widetilde{p}_{\widetilde{\mathbf{T}}_{c_{2},v_{4}}}^{4}\right)=f\left(\gamma_{2}^{0}+\gamma_{1}^{0}\gamma_{3}^{0}+\gamma_{2}^{0}\right)=(1+2+1)/2=2$. Therefore, for $v_{4}$, the modification of $\mathbf{T}_{3,4}$ from 00 to $1$ is the best choice to eliminate the maximum number of $6$-cycles without creating new $4$-cycles.
1 for $i=0$ to $I_{\text{MP}}-1$ do
2       for $v\in\mathbb{V}$ do
             $p_{v}^{l}\leftarrow{\small\textsf{MsgPass}}(\mathcal{G}_{\text{U}},l);$// follow (9)-(12)
3
4      $\widetilde{\mathbf{T}}_{c,v}^{\ast}$leftarrow (16);
5       Update $\mathcal{G}_{\text{U}}$ by $\widetilde{\mathbf{T}}_{c,v}^{\ast}$;
6
[…]
Algorithm 1 Optimization on Unified Graph $\mathcal{G}_{\text{U}}$
Based on the aforementioned method, we use Remark 2 to design a good spreading strategy on $\mathcal{G}_{\text{U}}$.
Remark 2.
Given a target of eliminating $l$-cycles, when only one edge allocation is altered, (16) provides the globally optimal to minimize the number of $l$-cycles in $\mathcal{G}_{\text{U}}$, without introducing any new shorter cycles.
$\widetilde{\mathbf{T}}_{c,v}^{\ast}\triangleq\operatorname*{argmax}\limits_{\widetilde{\mathbf{T}}_{c,v}}$ $\underbrace{\left(f\left(p_{v}^{l}\right)-f\left(\widetilde{p}_{\widetilde{\mathbf{T}}_{c,v}}^{l}\right)\right)/2}_{\varepsilon_{v}^{c}}$ (16)
[…]
Proof.
The metric $\varepsilon_{v}^{c}$ in (16) essentially measures the number of $l$-cycles containing a specific VN $v$ that can be removed if reallocating the edge $e_{v}^{c}$. Namely, we need to demonstrate that ${\varepsilon_{v}^{c}}$ is also the number of $l$-cycles removed in the entire $\mathcal{G}_{\text{U}}$. First, assume that there exists a scenario where the total number of cycles eliminated in the whole graph is less than ${\varepsilon_{v}^{c}}$ (the proof follows similarly if it is ‘greater’). For example, this edge allocation creates a new $l$-cycle that does not involve $v$. It can be easily shown that this new cycle must include the edge ${e_{v}^{c}}$ since it is caused by this edge reallocation. Considering that an edge cannot exist independently of nodes, the edge ${e_{v}^{c}}$ must include $v$, meaning this new cycle also involves $v$. This contradicts our initial assumption. Hence, Remark 2 always holds. ∎
In Algorithm 1, we summarize the approach to optimize the spreading strategy on the unified graph $\mathcal{G}_{\text{U}}$, where $I_{\text{MP}}$ is the maximum number of attempts for running message-passing (as well as the maximum number of edges that can be adjusted). The operations from (9) to (12) are summarized in the function ${\small\textsf{MsgPass}}(\cdot)$. Let ${e_{v}^{c}}^{\ast}$ and ${\varepsilon_{v}^{c}}^{\ast}$ denote the edge and the number of reduced $l$-cycles, respectively, corresponding to the optimal $\widetilde{\mathbf{T}}_{c,v}^{\ast}$. As shown in (16), we adjust the edge ${e_{v}^{c}}^{\ast}$ with the allocation of $\widetilde{\mathbf{T}}_{c,v}^{\ast}$ that achieves the greatest reduction in ${\varepsilon_{v}^{c}}^{\ast}$ $l$-cycles without any increase of shorter cycles. Algorithm 1 is therefore a greedy algorithm to eliminate the residual cycles in $\mathcal{G}_{\text{U}}$ in which each step identifies and modifies the edge that brings the steepest decline in cycle count.
Figure 6: Threshold analysis comparing PBRL codes from [17], 5G-NR LDPC codes from [4], and ESRL codes. In the lower figure, solid lines represent $\rho=0$ without puncturing, while dashed lines represent $\rho\neq 0$ with puncturing.
IV-B Modified Coupled Reciprocal Channel Approximation
RCA [44] is a fast and accurate approach to estimating the asymptotic iterative decoding threshold, which determines the minimum signal-to-noise-ratio (SNR) required for reliable transmission and has been successfully applied to the performance analysis of protograph-based LDPC ensembles under the AWGN channel [17]. In RCA, the SNR, denoted as $s$, is propagated along each edge of a protograph. Unlike [44] and [17], we need to introduce the coupling length $L$ into the RCA analysis for ESRL codes. Specifically, rather than using the uncoupled protomatrix $\mathbf{B}$, we perform RCA on the coupled protomatrix $\mathbf{B}_{\text{SC}}$ to incorporate the effects of edge spreading and termination, i.e., including matrices $\mathbf{T}$ and $\mathbf{Q}$, thus enhancing the accuracy of the threshold estimation.
Notably, the structured irregularity at the boundaries of SC-LDPC codes leads to a wave-like convergence, where the accumulated SNR of each bit progressively decreases from the ends toward the middle and the minimum accumulated SNR rises as $L$ increases [45]. Considering the threshold is determined by the minimum SNR that satisfies $s>T$ for all VNs [17], where $T$ is the stopping value in RCA, a limited $L$ results in a higher estimated threshold. Hence, the optimal threshold of an SC-LDPC code is achieved as $L\!\rightarrow\!\infty$, aligning with the results of density evolution (DE) [33] (the gap to the limit vanishes as $L$ increases). However, as $L$ grows, the $\mathbf{B}_{\text{SC}}$ size also increases significantly, making the RCA processing time-consuming [35]. To address this issue, we present a modified coupled RCA tailored to ESRL codes to approach the threshold of $L\!\rightarrow\!\infty$ with a finite $L$ on $\mathbf{B}_{\text{SC}}$.
Let $R(s)=C^{-1}(1-C(s))$ be the reciprocal energy function, where $C(s)$ calculates the channel capacity under a given SNR (defined in (5) of  [17]). Denote the SNR message passed from the $v$-th VN to the $c$-th CN as $s_{v,c}$, similarly, let $s_{c,v}$ be the message passed from the $c$-th CN to the $v$-th VN. At the $t$-th iteration, the SNR messages are updated as
{$s_{v,c}^{(t)}&=s_{v,c}^{(0)}+\sum\limits_{c^{\prime}\in\mathbb{C}_{v}\backslash c}R\left(s_{c^{\prime},v}^{(t-1)}\right),\\ s_{c,v}^{(t)}&=\sum\limits_{v^{\prime}\in\mathbb{V}_{c}\backslash v}R\left(s_{v^{\prime},c}^{(t)}\right).\end{aligned}\right.$ (17)
The initialization $s_{v,c}^{(0)}$ is set to input channel LLR if transmitted, or to 00 if punctured. Hence, the overall reliability of each VN is computed as $q_{v}^{(t)}=s_{v,c}^{(0)}+\sum_{c\in\mathbb{C}_{v}}R(s_{c,v}^{(t)})$, and $\bm{q}^{(t)}$ thus forms a vector on $\mathbf{B}_{\text{SC}}$ across all $(L\cdot n+\omega\cdot m)$ VNs of $\mathbf{B}_{\text{SC}}$. In (18), we first take the average of $\bm{q}^{(t)}$ over the dimension of $L$ spatial positions to mitigate the influence of a limited $L$ on wave-like convergence [45]
$\bar{\bm{q}}^{(t)}=\frac{1}{L}\sum_{i=0}^{L-1}\bm{q}^{(t)}_{i\cdot n:(i+1)\cdot n-1},$ (18)
where $\bm{q}^{(t)}_{i\cdot n:(i+1)\cdot n-1}$ represents a segment of $\bm{q}^{(t)}$ at spatial position $i$, each of which has a length of $n$. Since the accumulated SNRs on the tail are lower relative to $\bar{\bm{q}}^{(t)}$, to make (18) more accurate, we generate a weighted sum $\widehat{\bm{q}}^{(t)}$ as
$\widehat{\bm{q}}^{(t)}=\frac{L\cdot n}{L\cdot n+\omega\cdot m}\cdot\bar{\bm{q}}^{(t)}+\frac{\omega\cdot m}{L\cdot n+\omega\cdot m}\cdot\zeta^{(t)}_{\text{tail}}.$ (19)
We define the tail offset as $\zeta_{\text{tail}}^{(t)}=\frac{1}{\omega\cdot m}\sum_{v=L\cdot n}^{L\cdot n+\omega\cdot m-1}q^{(t)}_{v}$. Finally, after $I_{\text{RCA}}$ iterations, the minimum SNR that ensures that every element of $\widehat{\bm{q}}^{(t)}$ exceeds $T$ is recognized as the decoding threshold. In Fig. 6, the ESRL code designed in Section VI enjoys better thresholds across a wide range of rates compared to 5G-NR [4]. Besides, the puncturing scheme introduced in Section V-A can reduce thresholds by $0.12$ dB at high rates for the ESRL code, which will be discussed later.
Figure 7: Design flow of the ESRL code.
IV-C Design Flow of ESRL Codes
In this section, we present a comprehensive design flow for the proposed ESRL code. In Fig. 7, the process begins with basic parameters: $\{m^{\prime},n^{\prime},m,n,\rho,\omega,L,g,Z,M,N\}$. Herein, $m^{\prime}\times n^{\prime}$ represents the size of $\mathbf{B}_{\text{HRC}}$, while other parameters have been previously introduced, e.g., $M=\left(m-\rho\right)\cdot\left(L+\omega\right)\cdot Z$ and $N=\left(n\cdot L+m\cdot\omega-\rho\cdot(L+\omega)\right)\cdot Z$ refer to (5). The core of the design flow is divided into three procedures:  1 Design of the HRC at a fixed high rate  2 Design of the incremental matrix extension (IME), which integrates the IRC and SSC  3 Circulant lifting to achieve a target code length and meet the desired girth $g$.
The HRC design process is outlined in Algorithm 2. In Step 1, under the ESRL code constraints from Section III-B, DE is used to determine the optimal distribution of $d_{v}$, denoted as the polynomial $p_{d_{v}}^{\ast}$. Step 2 employs the progressive edge growth (PEG) algorithm [46] and the approximate cycle extrinsic message degree (ACE) constraint [47] to maximize local girth and to ensure the graph connectivity of short cycles on $\mathbf{B}_{\text{HRC}}$, respectively, which are two classical methods for designing finite-length LDPC codes. Notably, threshold analysis on $\mathbf{B}_{\text{HRC}}$ is not part of our flow, aligning with the objective of ESRL codes in Section III-A to focus on constructing a better coupled protomatrix $\mathbf{B}_{\text{SC}}$. The next step is to use the unified graph $\mathcal{G}_{\text{U}}$ in Section IV-A and the modified coupled RCA in Section IV-B to optimize $\mathbf{B}_{\text{SC}}$. In Step 3, the matrix $\mathbf{T}$ is initially set to $\mathbf{0}$, indicating that all edges start in $\mathbf{B}_{0}$. The unified graph $\mathcal{G}_{\text{U}}$ is then employed to optimize the spreading strategy, reallocating each edge to minimize cycle count. In Step 4, $\mathbf{B}_{\text{SC}}$ is constructed using the code profile $\mathbf{B}$$\mathbf{T}$, and $\mathbf{Q}$, followed by modified coupled RCA to compute the threshold. Since $\mathbf{Q}$ has a fixed structure without additional design, it directly participates in the threshold analysis of $\mathbf{B}_{\text{SC}}$. Finally, after $I_{\text{HRC}}$ iterations, the optimal $\mathbf{B}_{\text{HRC}}^{\ast}$ and $\mathbf{T}^{\ast}$ with the lowest threshold are selected.
1 Step 1: Find the optimal $p_{d_{v}}^{*}$ by DE
2 Step 2: Generate $\mathbf{B}_{\text{HRC}}$ through PEG + ACE
3    $\mathbf{B}_{\text{HRC}}\leftarrow{\small\textsf{PEG}}(p_{d_{v}}^{\ast},m^{\prime},n^{\prime})$;
4    if ${\small\textsf{ACE}}(\mathbf{B}_{\text{HRC}})<\eta_{\text{ACE}}$ then
5          Repeat Step 2;
6 Step 3: Optimize over unified graph $\mathcal{G}_{\text{U}}$
7    $\mathbf{T}\leftarrow\mathbf{0}$; //initialize as $\mathbf{B}_{0}$
8    $\mathbf{T}\leftarrow\mathcal{G}_{\text{U}}(\mathbf{B}_{\text{HRC}},\mathbf{T})$; //refer to Algorithm 1
9 Step 4: Perform modified coupled RCA
10    $\mathbf{B}_{\text{SC}}\leftarrow{\small\textsf{EdgeSpread}}(\mathbf{B}_{\text{HRC}},\mathbf{T},\mathbf{Q},L)$; //as (4)
11    $\xi\leftarrow{\small\textsf{RCA}}(\mathbf{B}_{\text{SC}})$;
12    if $\xi<\xi^{\ast}$ then
13          $\xi^{\ast}\leftarrow\xi,\;\mathbf{B}_{\text{HRC}}^{\ast}\leftarrow\mathbf{B}_{\text{HRC}},\;\mathbf{T^{\ast}\leftarrow\mathbf{T}}$;
14 Repeat Step 2 to Step 4 $I_{\text{HRC}}$ times;
return $\mathbf{B}_{\text{HRC}}^{\ast}$ and $\mathbf{T}^{\ast}$;
Algorithm 2 Design of the HRC
The IME design process is summarized in Algorithm 3. Compared to the overall design of the HRC in Algorithm 2, the IME is constructed row-wise until the full matrix of size $m\times n$ is completed. In the ESRL code, we find that the performance of the growing IME matrix is highly sensitive to the previously designed sections, especially after edge spreading, i.e., overemphasizing the performance of a specific intermediate rate will lead to significant performance degradation in subsequent matrix extensions. To address this issue, instead of using a greedy approach to explore a local optimum for each newly added row as in [17], we start with a function called ${\small\textsf{GlobalWeightRef}}(\cdot)$ in Step 1, which essentially runs Algorithm 2 to design a reference $\mathbf{B}$ of size $m\times n$ directly and thus offers a preliminary IME example. Note that although this reference is not optimal, it can provide a rough weight guideline $\mu_{\text{IME}}$ for the IME design to avoid being trapped in local optima for intermediate rows. Similarly, Step 2 involves using the PEG and ACE algorithms for the newly added row. In Step 3, the unified graph $\mathcal{G}_{\text{U}}$ is generated to optimize the spreading strategy. It is noteworthy that when designing the $c$-th row of the IME, the edges and corresponding allocations of previous $c-1$ rows must remain unchanged to preserve the optimization results obtained before. Besides, certain special structures in ESRL codes, such as assigning the diagonal of the SSC to $\mathbf{B}_{0}$, need to be retained. Finally, we run the process $I_{\text{IME}}$-times for each row to search for optimal $\mathbf{B}^{\ast}$ and $\mathbf{T}^{\ast}$ with the lowest threshold, until completing the entire ESRL protomatrix design $\mathbf{B}_{\text{SC}}$.
1 Step 1: Initialization
2    $\mathbf{B}^{\ast}\leftarrow\mathbf{B}_{\text{HRC}}^{\ast}$;
3    //Avoid trapping into local optima
4    $\mu_{\text{IME}}\leftarrow{\small\textsf{GlobalWeightRef}}(\mathbf{B}_{\text{HRC}},m,n)$;
5 Step 2: Generate IME row by row
6    for $c=m^{\prime}$ to $m-1$ then
7 Step 2.1: Add a row through PEG + ACE
8          $\mathbf{B}\leftarrow\mathsf{PEG}(\mathbf{B}^{\ast},\mu_{\text{IME}})$;
9          if ${\small\textsf{ACE}}(\mathbf{B})<\eta_{\text{ACE}}$ then
10                Repeat Step 2.1;
11 Step 2.2: Optimize over unified graph $\mathcal{G}_{\text{U}}$
12          $\mathbf{T}\leftarrow{\small\textsf{AppendOneEmptyRow}}(\mathbf{T}^{\ast})$; //initial
13          //No changes in previous edges
14          $\mathbf{T}\leftarrow\mathcal{G}_{\text{U}}(\mathbf{B},\mathbf{T})$;
15 Step 2.3: Perform modified coupled RCA
16          $\mathbf{B}_{\text{SC}}\leftarrow{\small\textsf{EdgeSpread}}(\mathbf{B},\mathbf{T},\mathbf{Q},L)$; //as (4)
17          $\xi\leftarrow\mathsf{RCA}(\mathbf{B}_{\text{SC}})$;
18          if $\xi<\xi^{*}$ then
19                $\xi^{*}\leftarrow\xi,\;\mathbf{B}^{*}\leftarrow\mathbf{B},\;\mathbf{T^{*}\leftarrow\mathbf{T}}$;
20          Repeat Step 2.1 to Step 2.3 $I_{\text{IME}}$ times;
return $\mathbf{B}^{*}$ and $\mathbf{T}^{*}$;
Algorithm 3 Design of the IME
The last step in the ESRL design flow is to incorporate the circulant lifting to transform $\mathbf{B}_{\text{SC}}$ into $\mathbf{H}_{\text{SC}}$, eliminating residual short cycles to meet the desired girth $g$. We define a matrix $\mathbf{P}$ of size $m\times n$ to store cyclic-shifting values of each non-zero entry in $\mathbf{B}$, where $\mathbf{P}_{c,v}\in[0,Z-1]$. [^3] Note that since the ESRL code has no parallel edges, the matrix $\mathbf{P}$ can be converted to $\mathbf{B}$ by binarization, i.e., $\mathbf{P}$ inherently contains all the information of $\mathbf{B}$. In practice, $\mathbf{B}$ can be replaced by $\mathbf{P}$ in the ESRL code profile, and we will no longer distinguish these two matrices in the remainder of this paper. According to the Fossorier condition [19], an $l$-cycle existing in $\mathcal{G}_{\text{U}}$ persists in $\mathbf{H}_{\text{SC}}$ only if the lifting path $\varphi$ of this cycle is also zero, as shown below [^4] For brevity, we skip the description of lifting on the tail matrix $\mathbf{Q}$ and the corresponding part in $\mathbf{H}_{\text{SC}}$, as the lifting process is similar.
$\varphi=\Delta\mathbf{P}_{0}+\Delta\mathbf{P}_{1}+\Delta\mathbf{P}_{2}+\ldots+\Delta\mathbf{P}_{l/2-1},$ (20)
where $\Delta\mathbf{P}_{i}=\mathbf{P}_{c_{i},v_{i}}-\mathbf{P}_{c_{i},v_{\text{mod}(i+1,l/2)}}$. Notably, (8) and (20) include an identical structure, which means that the circulant lifting step can also be incorporated into the optimization over $\mathcal{G}_{\text{U}}$. A design example will be provided in Section VI.
V Puncturing Scheme and Decoding Algorithm
In this section, we introduce the puncturing scheme and high-throughput SLME decoding tailored to ESRL codes.
V-A Puncturing Scheme
Punctured bits can significantly reduce the decoding threshold and support higher transmission rates. As shown in Fig. 6, both the PBRL [17] and 5G-NR LDPC codes [4] benefit substantially from punctured bits. In this section, we therefore also include puncturing in our ESRL codes.
As depicted in Fig. 8, the first VN at each spatial position of $\mathbf{B}_{\text{SC}}$ in the ESRL code is punctured, and these nodes are designed to connect with all CNs in the uncoupled protomatrix $\mathbf{B}$ (described in Section III-B). Note that the first column of the tail matrix $\mathbf{Q}$ is also punctured to maintain a regular structure. This puncturing scheme improves the ESRL decoding threshold by $0.12$ dB as visible in Fig. 6. Moreover, an efficient puncturing scheme should also ensure good decoding convergence within a limited number of iterations [32]. To evaluate this rapid recovery, we use the concept of a one-step-recoverable VN, as defined in [14]. Specifically, a punctured VN is called one-step-recoverable if it has at least one adjacent CN whose other VN neighbors are all unpunctured. Notably, the authors of [14] proved that such a punctured VN can be fully recovered after the first decoding iteration on the BEC. One-step-recoverable bits can also provide greater reliability on the AWGN channel compared to other types of punctured bits, leading to better performance with only a few iterations. This leads to Remark 3.
Figure 8: Puncturing scheme of the ESRL code, where the proposed semi-layered multi-engine decoding is running.
Remark 3.
All punctured VNs in the ESRL code belong to one-step-recoverable VNs, ensuring decoding convergence.
Proof.
Given that the protomatrix $\mathbf{B}$ is binary and contains no parallel edges, and only the first VN at each spatial position of $\mathbf{B}_{\text{SC}}$ is punctured, each CN thus connects to at most one punctured VN corresponding to an information bit. This is even stricter than the condition of one-step-recoverable VNs and Remark 3 always holds. Therefore, our puncturing scheme not only optimizes the decoding threshold with higher transmission rates but also ensures good decoding convergence within a limited number of iterations. ∎
V-B SLME Decoding of ESRL codes
In this section, we propose a high-throughput semi-layered decoding algorithm using a multi-engine architecture in [12], tailored to ESRL codes. This algorithm can be flexibly configured as either high-performance BP decoding of the entire $\mathbf{H}_{\text{SC}}$ or as low-complexity windowed BP sliding along the diagonal of $\mathbf{H}_{\text{SC}}$. Let $W$ denote the window size, $S$ the sliding step, and $I_{\text{max}}$ the maximum number of decoding iterations within the window at each step. For instance, by setting $W\!=\!S\!=\!L+\omega$, SLME decoding corresponds to traditional BP decoding with $I_{\text{max}}$ for ESRL codes.
Figure 9: Protomatrix example of ESRL codes is designed with $k=22$ and $\omega=1$, supporting a range of rates from $0.87$ to $0.34$ when $L=10$.
Figure 10: ACE spectrum of the protograph example of ESRL codes.
While coupling generally limits parallel decoding, ESRL codes without parallel edges can perform semi-layered decoding through a more compact multi-engine architecture compared to [12], achieving higher parallelism. Specifically, we apply individual layered decoding to each $\mathbf{U}$ (defined in (2) of Section III-A and also shown in Fig. 3), without causing memory read-write conflicts. In Fig. 8, with a window size of $W=4$, conventional layered decoding proceeds row-wise from top to bottom (indicated by a grey arrow). In contrast, SLME decoding can deploy four identical sub-decoders working in parallel (each running layered decoding shown by a light red arrow in Fig. 8), thus resulting in a 4×4$\times$4 × throughput gain in the example.
It is noteworthy that since posterior LLRs are updated across multiple rows simultaneously in multi-engine, it highlights the importance of static decoding schedule optimizations for ESRL codes. Inspired by [22], we introduce a simple static schedule optimization that prioritizes rows with smaller $d_{c}$, while postponing those with higher degrees to later in the sequence. This priority ensures that reliable messages propagate to more VNs. We define a layer ordering $\mathbb{O}$ as
$\mathbb{O}\triangleq\{o_{i},0\leq i<m|d_{c_{o_{0}}}\leq d_{c_{o_{1}}}\leq\ldots\leq d_{c_{o_{m-1}}}\}.$ (21)
The following simulation results for ESRL codes are based on this ordering $\mathbb{O}$. As mentioned before, setting $W=S=L+\omega$ transforms SLME decoding into traditional BP decoding on $\mathbf{H}_{\text{SC}}$. This configuration allows for a fair comparison with 5G-NR LDPC codes under equivalent computational complexity, which is discussed in Section VI.
VI Design Example and Performance Comparisons
In this section, we provide a design example of ESRL codes and conduct comprehensive comparisons with 5G-NR.
VI-A Design Example of ESRL Codes
Figure 11: FER performance of ESRL codes with different lifting sizes.
Fig. 9 illustrates an example ESRL protomatrix family designed for $k=22$ (in each batch). We intentionally set $k$ equal to that of 5G-NR to facilitate fair comparisons in Section VI-B. Following SC-LDPC decoder implementations [41], we set $\omega=1$ to simplify the code design. Moreover, the coupling length $L$ is set as $10$ to balance the rate loss and the underlying blocklength for a given frame length. Let $m^{\prime}=4$$n^{\prime}=26$$m=40$, and $n=62$. Hence, the HRC protomatrix has dimensions of $4\times 26$. The IME consists of the IRC protomatrix (sized $36\times 26$) and the SSC protomatrix (sized $36\times 36$). Considering that the first VN in the HRC is always punctured ($\rho=1$), the example ESRL code can support $37$ distinct rates, ranging from the highest rate of $0.87$ to the lowest rate of $0.34$. The uncoupled protomatrix $\mathbf{B}$ has a size of $40\times 62$ and a weight of $314$. Note that with this configuration we mark six classical rates in Fig. 9: $0.87~{}(4\times 26)$$0.8~{}(6\times 28)$$0.72~{}(9\times 31)$$0.63~{}(13\times 35)$$0.52~{}(20\times 42)$, and $0.34~{}(40\times 62)$, which are also supported by 5G-NR and are used for the subsequent comparisons. For instance, the highest rate is computed as $R=(22\times 10)/(26\times 10+4-1\times(10+1))\approx 0.87$ using (5). The edge-spreading matrix $\mathbf{T}$ is also reported in Fig. 9, with $\mathbf{B}_{0}$ and $\mathbf{B}_{1}$ distinguished by colors. As mentioned in Section III-C, the tail matrix $\mathbf{Q}$ has a fixed tri-diagonal structure, marked as black in Fig. 9, since it does not participate in the spreading process. In terms of connectivity of short cycles, in the lowest rate protograph, every cycle of length up to $8$ is connected to at least $12$ extrinsic CNs. For $4$-cycles, the value of ACE $\eta$ ranges from $6$ to $12$. Fig. 10 also shows the ACE spectrum of $6$-cycles and $8$-cycles. It is noteworthy that given these ACE values are obtained from the protograph before lifting, the actual graph connectivity will be even better in $\mathbf{H}_{\text{SC}}$ of the ESRL code after lifting.
Fig. 11 shows the frame error rate (FER) performance of the example ESRL codes under different lifting sizes. Let $\mathcal{C}$ define a series of rate-compatible LDPC codes with $K$ information bits, and $\mathcal{C}(N,K)$ be one of its code patterns with a rate $R=K/N$. Based on $Z\in\{39,78,390\}$, three series of ESRL codes are defined: $\mathcal{C}_{\text{SC}}^{\text{I}}$ $(K=8580)$$\mathcal{C}_{\text{SC}}^{\text{II}}$ $(K=17160)$, and $\mathcal{C}_{\text{SC}}^{\text{III}}$ $(K=85800)$, where $K\!=\!k\cdot L\cdot Z$. The maximum frame lengths of $\mathcal{C}_{\text{SC}}^{\text{I}}$$\mathcal{C}_{\text{SC}}^{\text{II}}$, and $\mathcal{C}_{\text{SC}}^{\text{III}}$ are $N=25311$$N=50622$, and $N=253110$ for the lowest rate of $R=0.34$, respectively, which are approximately 1×1$\times$1 ×2×2$\times$2 ×, and 10×10$\times$10 × longer compared to 5G-NR LDPC codes with the same rate. [^5] Note that in this paper, we only consider the maximum blocklength 5G-NR LDPC codes with $Z=384$ to fully present the 5G-NR performance. In SLME decoding, we let $W=S=L+\omega=11$, thus decoding is configured to run BP on the entire $\mathbf{H}_{\text{SC}}$ of ESRL codes. Given that hardware implementations as in [22] and [48] typically limit the iteration count to $5$ or below to minimize decoding delay and enhance throughput, in Fig. 11, we set $I_{\text{max}}=5$ to highlight the ESRL performance in these high-throughput scenarios. At $R=0.87$, since the ESRL code only uses the HRC part without involving the IME, a slight error floor appears at the FER of $10^{-5}$. At other rates, ESRL codes exhibit a FER performance without error floors and achieve stable performance gains when $Z$ increases. For instance, $\mathcal{C}_{\text{SC}}^{\text{III}}$ with $Z=390$ outperforms $\mathcal{C}_{\text{SC}}^{\text{I}}$ with $Z=39$ by $0.35$ dB when $R=0.72$.
VI-B Comparisons with 5G-NR LDPC Codes
Figure 12: FER comparisons of ESRL codes and 5G-NR LDPC codes [4] (including their variants [49, 33, 34, 35]) using L-NMS decoding with $I_{\text{max}}=5$.
Comparing the performance of ESRL codes and 5G-NR LDPC codes in a reasonably fair manner involves many factors. As stated in [3], computational complexity does not always accurately reflect the true silicon area cost, and throughput is significantly affected by the hardware architecture and the level of processing parallelism. Therefore, in addition to using the same code configurations, all decoding parameters and complexity should be the same in both software and hardware as much as possible.
To concentrate on the selection of channel coding without the influence of decoding algorithms and hardware architectures, we adopt the following approach. Specifically, we intentionally extend the uncoupled blocklength of ESRL codes to match that of 5G-NR LDPC codes. Combined with the multi-engine architecture in Section V-B, when SLME is configured as BP decoding on $\mathbf{H}_{\text{SC}}$, the corresponding ESRL decoder can be decomposed into $L$ sub-decoders operating in parallel (with the tail redundancy causing some extra overhead). Each sub-decoder is similar in size to that of a single 5G-NR LDPC decoder [22]. In this context, compared to multi-core 5G-NR LDPC decoders [50], we can run the same decoding algorithm with the same hardware implementation for ESRL codes to eliminate additional impacts.
For the comparisons, we select the ESRL code $\mathcal{C}_{\text{SC}}^{\text{III}}$, where each batch contains $8580$ information bits, and the entire frame consists of a total of $85800$ bits. Meanwhile, we define $\mathcal{C}_{\text{5G}}^{\text{I}}$ as a packet frame that contains $10$ individual 5G-NR LDPC codewords, with each code block containing $8448$ information bits. We employ a multi-core architecture to decode $\mathcal{C}_{\text{5G}}^{\text{I}}$, i.e., each sub-decoder independently works on an individual 5G-NR LDPC block. It is noteworthy that each block in both $\mathcal{C}_{\text{SC}}^{\text{III}}$ and $\mathcal{C}_{\text{5G}}^{\text{I}}$ has a similar size, and their overall frame lengths are also comparable. For each block in both the multi-engine architecture used for ESRL codes and the multi-core setup for 5G-NR LDPC codes, we employ the classical layered normalized min-sum (L-NMS) decoding and adopt the state-of-the-art block-parallel 5G-NR LDPC decoder in [22].
Figure 13: $E_{b}/N_{0}$ at the FER of $10^{-5}$ against $I_{\text{max}}$.
Figure 14: FER comparisons of ESRL codes and 5G-NR LDPC codes using L-SP decoding with $I_{\text{max}}=200$.
For high-throughput scenarios with $I_{\text{max}}=5$, as shown in Fig. 12, ESRL codes exhibit significant performance advantages compared to 5G-NR LDPC codes across various rates. For instance, at $R=0.8$, the ESRL code outperforms the 5G-NR LDPC code by $1.15$ dB, and although the gain slightly decreases it remains $0.65$ dB at $R=0.34$. Note that all algorithmic complexity and implementation complexity have been normalized, these gains are only attributable to channel coding. Moreover, to fully present the FER performance of 5G-NR LDPC codes, we refer to [49] and define such extended 5G-NR LDPC codes as $\mathcal{C}_{\text{5G}}^{\text{II}}$ $(K=84480)$ by redesigning a 5G-NR LDPC exponent matrix with a lifting size 10×10$\times$10 × larger $(Z=3840)$ and a non-decreasing girth. The FER of $\mathcal{C}_{\text{5G}}^{\text{II}}$, referred to as ‘5G-NR-Plus’, is plotted in Fig. 12. Given that the blocklength of the original 5G-NR LDPC codes is already on the order of tens of thousands, a 10×10$\times$10 × blocklength increase can bring only a $0.1$ dB gain compared to the original 5G-NR LDPC codes. Hence, ESRL codes still far outperform 5G-NR-Plus LDPC codes [49]. Subsequently, to intuitively highlight the difference between ESRL codes and previous protograph-based rate-compatible SC-LDPC code construction in [33, 34, 35], we replace the uncoupled $\mathbf{B}$ of ESRL codes with the 5G-NR matrix and define the code $\mathcal{C}_{\text{5G}}^{\text{III}}$. All other ESRL optimizations, such as the unified graph $\mathcal{G}_{\text{U}}$ and the tail matrix $\mathbf{Q}$, are still preserved in $\mathcal{C}_{\text{5G}}^{\text{III}}$ to ensure a fair comparison. This more classical SC-LDPC construction is referred to as ‘5G-NR-Coupled’ in Fig. 12. When taking $R=0.63$ as an example, edge spreading for the 5G-NR matrix brings a $0.4$ dB gain from coupling at the FER of $10^{-5}$, yet, compared to the ESRL code, the remaining $0.55$ dB gain is due to the well-designed ESRL code matrix. This demonstrates the significance of designing the optimal coupled matrix in ESRL codes, adhering to the design motivation behind ESRL codes.
Fig. 13 illustrates the relationship between $E_{b}/N_{0}$ at the FER of $10^{-5}$ against $I_{\text{max}}$ for ESRL codes and 5G-NR LDPC codes. At $R=0.52$, to match the ESRL performance of $I_{\text{max}}=5$, the 5G-NR LDPC code needs to enhance $I_{\text{max}}$ to $7$. This gap widens further when $R=0.8$. We summarize all comparisons of ESRL codes and 5G-NR LDPC codes in Table I. As shown in Table I, when $I_{\text{max}}=5$, the ESRL decoder can deliver a peak throughput of $209.4$ Gbps, while the 5G-NR LDPC decoder achieves a peak throughput of $195.4$ Gbps. These results are calculated by (15) of [22]. Notably, ESRL codes offer gains of more than $0.8$ dB at most rates. For 5G-NR with $I_{\text{max}}=7$, although the FER performance becomes comparable, the peak throughput drops to $139.5$ Gbps, indicating a $50.1\%$ throughput improvement for ESRL codes. For high-reliability scenarios, we use layered sum-product (L-SP) decoding [7] to replace L-NMS. Under $I_{\text{max}}=200$, ESRL codes outperform 5G-NR LDPC codes, especially at high code rates, e.g., this gain is $0.36$ dB at the FER of $10^{-5}$ when $R=0.87$. This greatly emphasizes the potential of ESRL codes for next-generation channel coding.
TABLE I: Comparisons between ESRL codes and 5G-NR LDPC codes.
  ESRL 5G-NR
  Max. Framelength $253110$ $253440$
Max. Blocklength $23790$ $25344$
Block Inf. Length $8580$ $8448$
Architecture Multi-Engine Multi-Core
Decoder Cores $10$ $10$
Algorithm SLME (L-NMS) L-NMS
Area 10$\approx$ 10$\times$ 10 × 10×10$\times$10 ×
Routing $Z=390$ $Z=384$
  Complexity $I_{\text{max}}=5$ $I_{\text{max}}=5$ $I_{\text{max}}=7$
Peak T/P [Gbps] $209.4$ $195.4$ $139.5$
FER@$10^{-5}$ $0.95$ dB Better Baseline $\approx$ ESRL
    Estimated by the posterior LLR memory overhead in hardware.
    Measured by the actual number of iterations running at each CN.
    Peak throughout is attained with $R=0.87$.
    Comparison is conducted at the FER of $10^{-5}$ at $R=0.52$.
VII Conclusion
In this work, we propose a new SC-LDPC code family called ESRL LDPC codes, which concentrates on constructing an optimal matrix after coupling, resulting in outstanding performance while maintaining extensive rate compatibility. Compared to LDPC block codes, ESRL codes demonstrate superior performance, thanks to the convolutional structure of the coupled matrix, as well as high throughput potential due to the regular and compact structure of underlying block codes.
Beyond the structure, properties, and code profile of ESRL codes, we also present the complete design flow and optimization strategies, including a unified graph to optimize the edge-spreading process and a modified coupled RCA to analyze the decoding threshold. In addition, we propose a high-throughput decoding algorithm called SLME for ESRL codes, which can be decomposed into multiple sub-decoders running in parallel. It is noteworthy that we provide a design example of ESRL codes and conduct comprehensive comparisons with 5G-NR LDPC codes and their variants, such as extended or coupled 5G-NR LDPC codes. When all decoding parameters and complexity are the same, ESRL codes significantly outperform 5G-NR LDPC codes in terms of error rate and throughput, making them a promising solution for 6G wireless systems.